Eagle Scout Court of Honor

for David Fifield July 22, 2001, 2:00–5:00 pm

Opening Color guard Invocation

Introduction Introduction of Eagle candidate Explanation of rank and requirements Description of service project

Eagle Charge Recognition of Eagles Eagle charge

Presentation Presentation of award Parents’ recognition

Closing Eagle Scout’s remarks Color guard Reception

“Alas, that Spring should vanish with the Rose! That Youth’s sweet-scented Manuscript should close!

The Nightingale that in the Branches sang, Ah, whence, and whither flown again, who knows!”

— From the Rubáiyát